OpenStudy (anonymous):

Evaluate. Please show all of your work. -y – (-2y – 4), for y = -3

OpenStudy (anonymous):

lets plug in y=-3 for y so we got -(-3) - ((-2)(-3) -4) first you need to know if you are multiplying by negative it becomes a positive so by applying BEDMAS let us do brackets first we get 3 - (6 - 4) = 3 -2 = 1 :)

OpenStudy (anonymous):

thanks